# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 1.1. Relatively big projectives
We introduce the notion of relatively big projective modules with respect to an ideal. This definition was first introduced by P. Příhoda in [P2] as a way of generalizing Bass’ uniformly big projective modules. We motivate the definition of this class of projective modules by comparing it with a characterization of the class of modules generated by a projective module.
Definition 1.10.
Let $\Lambda$ be a ring, and let $P$ be a projective right $\Lambda$-module. We define the class $\mathrm{Gen}(P)$ of modules generated by $P$ as the class of right $\Lambda$-modules which are a homomorphic image of a direct sum of copies of $P$.
@@
Let $\Lambda$ be a ring, let $P$ be a countably generated projective right $\Lambda$-module, and let $J$ be an ideal of $\Lambda$. The module $P$ is said to be big with respect to $J$, or $J$-big, if for every countably generated projective right $\Lambda$-module $Q$ satisfying $QJ=Q$, then $Q$ is a homomorphic image of $P$.
If $P$ is $J$-big projective and $J$ does not contain a nonzero element of $\mathcal{T}(\Lambda)$, then Definition 1.12 is only fulfilled for the trivial case $Q=\{0\}$. If $0\neq K\in\mathcal{T}(\Lambda)$ is such that $K\subseteq J$, then $P$ is $K$-big.
We summarize what it means to be relatively big with respect to an ideal which is the trace of a countably generated projective module.
@@
Let $\Lambda$ be a ring, and let $P,Q$ be countably generated projective right $\Lambda$-modules with trace ideals $I$ and $J$, respectively. Then
@@
    If $P$ is $J$-big such that $P/PJ$ is finitely generated, then $J$ is the minimal ideal of $\Lambda$ such that $P/PJ$ is finitely generated.
@@
Conversely, if $P\oplus Q^{(\omega)}\cong P$, then $Q^{(\omega)}$ is isomorphic to a direct summand of $P$, so there exists an epimorphism $P\to Q^{(\omega)}$. By Lemma 1.5, $QJ=Q$, so $P$ is $J$-big.
@@
[wiegand, Proposition 3.4] Let $\Lambda$ be a ring, let $J\in\mathcal{T}(\Lambda)$, and let $P$ be a $J$-big projective right $\Lambda$-module. Then the isomorphism class of $P$ is determined by the ideal $J$ and the isomorphism class of $P/PJ$, i.e., if $P^{\prime}$ is a $J$-big projective right $\Lambda$-module and $P^{\prime}/P^{\prime}J\cong P/PJ$ then $P^{\prime}\cong P$.
1.2. Finite generation modulo an ideal and the ()(*)( ∗ )-condition
@@
We recall the following crucial properties of a descending sequence of right ideals associated to a column finite idempotent matrix. The result is taken from [traces], but all the needed ideas are already in [P2].
@@
[traces, Lemmas 5.1 and 5.2] Let $\Lambda$ be a ring. Let $A=(a_{i,j})$ be a countable column finite matrix with entries in $\Lambda$ such that $A^{2}=A$, and set $P=A\Lambda^{(\mathbb{N})}$. For any $k\geq 0$, let $I_{k}=\sum_{i>k,j\in\mathbb{N}}a_{ij}\Lambda$. Then,
@@
    $\mathcal{I}(P)$ has minimal elements if and only if there exists $k_{0}\in\mathbb{N}$ such that, for any $\ell\in\mathbb{N}$, $\Lambda I_{k_{0}}=\Lambda I_{k_{0}+\ell}$. In this case $I_{k_{0}}=I_{k_{0}}^{2}$ and $\Lambda I_{k_{0}}$ is the minimal element of $\mathcal{I}(P)$.
@@
[P2] A ring $\Lambda$ satisfies condition ()(*)( ∗ ) provided every descending chain of two-sided ideals
@@
The following statement explains the role of condition ()(*)( ∗ ) in our discussion.
@@
Let $\Lambda$ be a ring. Let $P$ be a countably generated projective right $\Lambda$-module. If $\Lambda$ satisfies condition ()(*)( ∗ ), then $\mathcal{I}(P)$ has a minimal element and this minimal element is an idempotent ideal.
Proof.  By Lemma 1.15 $(v)$, $\mathcal{I}(P)$ has a minimal element if and only if the descending sequence of two-sided ideals $(\Lambda I_{k})_{k\geq 0}$ is stationary. By Lemma 1.15 $(iii)$, this will happen if and only if the partial sub-sequence $(J_{n})_{n\geq 0}$ given by that statement is stationary. Since such partial sub-sequence satisfies that $J_{n+1}J_{n}=J_{n+1}$, for any $n\geq 0$, it is stationary because $\Lambda$ satisfies condition ()(*)( ∗ ).
@@
[P2, Lemma 2.8] Let $\Lambda$ be a noetherian ring. Let $P$ be a countably generated projective right $\Lambda$-module such that $\mathcal{I}(P)$ has a minimal element $J$. Then $J$ is the trace ideal of a countably generated projective right $\Lambda$-module, and $P$ is $J$-big. In particular, if $\Lambda$ satisfies condition ()(*)( ∗ ), then for any countably generated projective right module $P$, there exists an ideal $J$ such that:
  • (i)
    $J$ is the trace of a countably generated projective right $\Lambda$-module,
  • (ii)
    $P$ is $J$-big, and
  • (iii)
    $P/PJ$ is finitely generated.
@@
$(2)$ We do not know whether condition ()(*)( ∗ ) is left-right symmetric, even for the case of noetherian rings.
$(3)$ There are noetherian rings with infinite descending sequences of idempotent ideals, so they do not satisfy condition ()(*)( ∗ ). Over such rings one constructs a countably generated projective module $P$ such that $\mathcal{I}(P)$ does not have a minimal element. See, for example, the last section of [traces].
@@ 2. The monoid of relatively big projective modules
We want to focus on $J$-big projective modules $P$ such that $P/PJ$ is finitely generated. This motivates the following definition.
@@
Let $\Lambda$ be a ring. We denote by $V(\Lambda)\sqcup B(\Lambda)$ the subset of $V^{*}(\Lambda)$ of isomorphism classes of countably generated projective right $\Lambda$-modules $P$ such that there exists a countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ satisfying:
  1. (i)
    $P$ is $J$-big, that is, $P\oplus Q^{(\omega)}\cong P$.
  2. (ii)
    $P/PJ$ is a finitely generated projective right $\Lambda/J$-module.
Remark 2.2.
Let $\Lambda$ be a ring. Then, $B(\Lambda)$ is the subset of $V^{*}(\Lambda)\setminus V(\Lambda)$ of isomorphism classes of countably generated projective right $\Lambda$-modules $P$, which are not finitely generated, and such that there exists a non-zero countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ satisfying $(i)$ and $(ii)$ in the previous definition.
Note that a countably generated projective $\Lambda$-module whose isomorphism class is in $B(\Lambda)$ is never indecomposable.
Example 2.3.
Let $\Lambda$ be a ring, let $P$ be a non-zero countably generated projective right $\Lambda$-module, and let $X$ be a finitely generated projective right $\Lambda$-module. Then
  1. (i)
    $[X]\in V(\Lambda)\sqcup B(\Lambda)$ with $Q=\{0\}$ and $J=\{0\}$.
  2. (ii)
    $[X\oplus P^{(\omega)}]\in B(\Lambda)$ with $Q=P$ and $J=\operatorname{Tr}_{\Lambda}(P)$.
@@
Proof.  $(i)$. Let $[P_{1}],[P_{2}]\in B(\Lambda)$. Then, for each $i=1,2$, there exists a countably generated projective right $\Lambda$-module $Q_{i}$ with trace ideal $J_{i}$ such that $P_{i}\oplus Q_{i}^{(\omega)}\cong P_{i}$ and $P_{i}/P_{i}J_{i}$ is a finitely generated projective right $\Lambda/J_{i}$-module. Take $P=P_{1}\oplus P_{2}$ and $Q=Q_{1}\oplus Q_{2}$. Then
@@
On the other hand, by Lemma 1.5, $J=\operatorname{Tr}_{\Lambda}(Q)=\operatorname{Tr}_{\Lambda}(Q_{1})+\operatorname{Tr}_{\Lambda}(Q_{2})=J_{1}+J_{2}$, so
@@
and $P/PJ$ is a finitely generated projective right $\Lambda/J$-module because it is a direct sum of quotients of finitely generated projective modules. Therefore $[P]\in B(\Lambda)$.
$(ii)$ is proved similarly.
@@
Let $\Lambda$ be a ring, and let $I\in\mathcal{T}(\Lambda)$. Then there is a semigroup embedding $V(\Lambda/I)\hookrightarrow V(\Lambda)\sqcup B(\Lambda)$. [^1] If $I\neq 0$, it cannot be made into a monoid embedding because $[0]$ maps to the isomorphism class of a non-finitely generated projective module.
Proof.  Let $Q$ be a countably generated projective right $\Lambda$-module with trace ideal $I$. Let $P^{\prime}$ be a projective right $\Lambda/I$-module. By Lemma 1.9, there exists a projective right $\Lambda$-module $P$ such that $I\subseteq\operatorname{Tr}_{\Lambda}(P)$, $P/PI\cong P^{\prime}$ and $\operatorname{Tr}_{\Lambda}(P)/I=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
Note that $P\oplus Q^{(\omega)}$ is $I$-big, since $P\oplus Q^{(\omega)}\cong P\oplus Q^{(\omega)}\oplus Q^{(\omega)}$. By Lemma 1.5, $QI=Q$, so $(P\oplus Q^{(\omega)})/(P\oplus Q^{(\omega)})I\cong P/PI$ is finitely generated. Therefore, $[P\oplus Q^{(\omega)}]\in V(\Lambda)\sqcup B(\Lambda)$. The isomorphism class of $P\oplus Q^{(\omega)}$ is independent on the choice of $Q$: let $Q^{\prime}$ be another countably generated projective right $\Lambda$-module with trace ideal $I$. Then, by Lemma 1.5, $Q^{\prime}I=Q^{\prime}$, so $(P\oplus Q^{\prime(\omega)})/(P\oplus Q^{\prime(\omega)})I\cong P/PI$. Hence, by Lemma 1.14, $[P\oplus Q^{(\omega)}]=[P\oplus Q^{\prime(\omega)}]$.
Let $\iota_{I}\colon V(\Lambda/I)\longrightarrow V(\Lambda)\sqcup B(\Lambda)$ be the map defined by $\iota_{I}[P^{\prime}]=[P\oplus Q^{(\omega)}]$. It is easy to check that this is well-defined: if $[P_{1}^{\prime}]=[P_{2}^{\prime}]$, then $P_{1}/P_{1}I\cong P_{2}/P_{2}I$. Since $P_{i}\oplus Q^{(\omega)}$ is $I$-big and $(P_{i}\oplus Q^{(\omega)})/(P_{i}\oplus Q^{(\omega)})I\cong P_{i}/P_{i}I$ for each $i=1,2$, by Lemma 1.14, $[P_{1}\oplus Q^{(\omega)}]=[P_{2}\oplus Q^{(\omega)}]$ and then $\iota_{I}[P_{1}^{\prime}]=\iota_{I}[P_{2}^{\prime}]$. It is easy to see that it is a semigroup homomorphism:
$\iota_{I}[P_{1}^{\prime}\oplus P_{2}^{\prime}]$ $=[P_{1}\oplus P_{2}\oplus Q^{(\omega)}]=[P_{1}\oplus Q^{(\omega)}\oplus P_{2}\oplus Q^{(\omega)}]$
$=[P_{1}\oplus Q^{(\omega)}]+[P_{2}\oplus Q^{(\omega)}]=\iota_{I}[P_{1}^{\prime}]+\iota_{I}[P_{2}^{\prime}].$
Finally, to see that it is injective, let $[P_{1}^{\prime}],[P_{2}^{\prime}]\in V(\Lambda/I)$ such that $\iota_{I}[P_{1}^{\prime}]=\iota_{I}[P_{2}^{\prime}]$. Then $[P_{1}\oplus Q^{(\omega)}]=[P_{2}\oplus Q^{(\omega)}]$, so $P_{1}\oplus Q^{(\omega)}\cong P_{2}\oplus Q^{(\omega)}$. By Lemma 1.5, $QI=Q$, so $(P_{i}\oplus Q^{(\omega)})/(P_{i}\oplus Q^{(\omega)})I\cong P_{i}/P_{i}I\cong P_{i}^{\prime}$ for each $i=1,2$, and hence $[P_{1}^{\prime}]=[P_{2}^{\prime}]$.
@@
Consider the set $\mathcal{T}(\Lambda)$ and the disjoint union $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$. We will denote the elements of $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$ by $([P],I)$ with $[P]\in V(\Lambda/I)$. Note that, if $I\subseteq J$, there is a morphism of monoids $f_{IJ}\colon V(\Lambda/I)\to V(\Lambda/J)$ defined by $f_{IJ}([X],I)=([X/XJ],J)$ with $[X]\in V(\Lambda/I)$. Therefore, $f_{IJ}(V(\Lambda/I))\subseteq V(\Lambda/J)$. Define addition in $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$ by
$([P_{1}],I)+([P_{2}],J)$ $=f_{I,I+J}([P_{1}],I)+f_{J,I+J}([P_{2}],J)$
@@
for each $[P_{1}]\in V(\Lambda/I)$, $[P_{2}]\in V(\Lambda/J)$. Therefore, the universal property of the disjoint union and Lemma 2.5 implies that there exists a semigroup embedding
@@
The following lemma states that this embedding is precisely an isomorphism.

### v2

@@ 1.1. Relatively big projectives
First we introduce the auxiliary notion of big projective module with respect to an ideal. It was first introduced by P. Příhoda in [P2] as a way of generalizing Bass’ uniformly big projective modules. We motivate the definition of this class of projective modules by comparing it with a characterization of the class of modules generated by a projective module.
Let $\Lambda$ be a ring, and let $M$ be a right $\Lambda$-module. We denote by $\mathrm{Gen}(M)$ the class of modules generated by $M$, that is the class of right $\Lambda$-modules which are a homomorphic image of a direct sum of copies of $M$.
@@
Let $\Lambda$ be a ring, let $P$ be a countably generated projective right $\Lambda$-module, and let $J$ be an ideal of $\Lambda$. The module $P$ is said to be big with respect to $J$, or $J$-big, if every countably generated projective right $\Lambda$-module $Q$ satisfying $QJ=Q$ is a homomorphic image of $P$.
The following remarks show that the significant case of Definition 1.11 is when $J$ is a trace ideal of a projective module.
Remarks 1.12.
Let $\Lambda$ be a ring, and let $J$ be an ideal of $\Lambda$.
  1. (1)
    If $Q$ is a countably generated projective right $\Lambda$-module such that $QJ=Q$ then $\operatorname{Tr}_{\Lambda}(Q)\leq J$. Therefore, if $J$ does not contain a non-zero element of $\mathcal{T}(\Lambda)$ then, by Lemma 1.5, Definition 1.11 is trivially fulfilled, so any countably generated projective module is $J$-big or, equivalently, 0-big.
  2. (2)
    If $P$ is $J$-big, and $K$ is a non-zero element of $\mathcal{T}(\Lambda)$ such that $K\subseteq J$, then $P$ is $K$-big (cf. Lemma 1.5).
    Observe that
    $J^{\prime}=\sum_{\begin{subarray}{c}K\leq J\\ K\in\mathcal{T}(\Lambda)\end{subarray}}K$
    is again a trace ideal of a projective module by Lemma 1.5. Then a countably generated projective module is $J$-big if and only if it is $J^{\prime}$-big.
    We stress that $J^{\prime}$ does not need to be the trace of a countably generated projective $\Lambda$-module. So it is not necessarily an element in $\mathcal{T}(\Lambda)$.
In the next lemma, we summarize what it means to be $J$-big with respect to $J\in\mathcal{T}(\Lambda)$.
@@
Let $\Lambda$ be a ring, and let $P,Q$ be countably generated projective right $\Lambda$-modules. Let $J=\operatorname{Tr}_{\Lambda}(Q)$. Then
@@
    If $P$ is $J$-big and $P/PJ$ is finitely generated, then $J$ is the minimal ideal of $\Lambda$ such that $P/PJ$ is finitely generated.
@@
Conversely, if $P\oplus Q^{(\omega)}\cong P$, then $Q^{(\omega)}$ is isomorphic to a direct summand of $P$, so there exists an epimorphism $P\to Q^{(\omega)}$. Any countably generated projective module with trace contained in $J$ is an epimorphic image of $Q^{(\omega)}$ by Lemma 1.10, so $P$ is $J$-big.
@@
The following lemma states that the isomorphism class of a $J$-big projective right $\Lambda$-module $P$ satisfying that $P/PJ$ is finitely generated, is determined, up to isomorphism, by the ideal $J$ and the isomorphism class of $P/PJ$.
@@
Let $\Lambda$ be a ring, let $J\in\mathcal{T}(\Lambda)$, and let $P$ be a $J$-big module such that $P/PJ$ is finitely generated. Then:
  • (i)
    If $P^{\prime}$ is a projective module isomorphic to $P$, then $P^{\prime}$ is $J$-big and $P^{\prime}/P^{\prime}J$ is finitely generated.
  • (ii)
@@
    [wiegand, Proposition 3.4] Let $P^{\prime}$ be a countably generated projective module that is $J^{\prime}$-big for some $J^{\prime}\in\mathcal{T}(\Lambda)$ and such that $P^{\prime}/P^{\prime}J^{\prime}$ is finitely generated. Then $P\cong P^{\prime}$ if and only if $J=J^{\prime}$ and $P/PJ\cong P^{\prime}/P^{\prime}J(=P^{\prime}/P^{\prime}J^{\prime})$.
Proof. We only need to prove $(i)$. Notice that a module isomorphic to a $J$-big module is also $J$-big. Moreover, if $P\cong P^{\prime}$ then $P^{\prime}/P^{\prime}J$ is a homomorphic image of $P/PJ$. So if $P/PJ$ is finitely generated, then so is $P^{\prime}/P^{\prime}J$.
Now we give the definition of relatively big projective which is crucial for the rest of the paper.
Definition 1.15.
Let $\Lambda$ be a ring, and let $P$ be a countably generated projective right $\Lambda$-module. The module $P$ is said to be relatively big if there exists a countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ satisfying:
  1. (1)
    $P$ is $J$-big, that is, $P\oplus Q^{(\omega)}\cong P$.
  2. (2)
    $P/PJ$ is a finitely generated projective right $\Lambda/J$-module.
We will also use the terminology $P$ is relatively $J$-big when we want to specify the ideal $J\in\mathcal{T}(\Lambda)$.
Example 1.16.
Let $\Lambda$ be a ring. Let $X$ be a finitely generated projective right $\Lambda$-module, and let $P$ be a countably generated projective right $\Lambda$-module. Then $X\oplus P^{(\omega)}$ is relatively big: take $Q=P$ and $J=\operatorname{Tr}_{\Lambda}(P)$ in the Definition 1.15.
In particular, taking $X=\{0\}$, we see that $P^{(\omega)}$ is relatively big for any countably generated projective module $P$. Also, taking $P=\{0\}$ we see that any finitely generated projective module is relatively big.
Example 1.16, is a particular instance of the following general construction of relatively big projective modules.
Lemma 1.17.
Let $\Lambda$ be a ring. Let $P$ be a countably generated projective right $\Lambda$-module, and let $J$ be the trace ideal of a countably generated projective right $\Lambda$-module $Q$. If $P/PJ$ is finitely generated, then $P\oplus Q^{(\omega)}$ is relatively $J$-big.
Proof. Note that $(P\oplus Q^{(\omega)})/(P\oplus Q^{(\omega)})J\cong P/PJ$, which is finitely generated. Since $P\oplus Q^{(\omega)}\cong P\oplus Q^{(\omega)}\oplus Q^{(\omega)}$, by Lemma 1.13, $P\oplus Q^{(\omega)}$ is $J$-big.
Remark 1.18.
Note that an indecomposable relatively big projective module must be finitely generated.
The following result characterizes when every relatively big projective module is isomorphic to a direct sum of finitely generated projective modules.
Proposition 1.19.
Let $\Lambda$ be a ring. Then every relatively big projective right $\Lambda$-module is a direct sum of finitely generated projective modules if and only if the following two properties are satisfied:
  1. (1)
    If $J$ is the trace ideal of a projective right $\Lambda$-module, then there is a right $\Lambda$-module $P$, which is a direct sum of finitely generated projective modules such that $\operatorname{Tr}_{\Lambda}(P)=J$.
  2. (2)
    If $J$ is the trace ideal of a countably generated projective right $\Lambda$-module and $P^{\prime}$ is a finitely generated projective right $\Lambda/J$-module, then there is a finitely generated projective right $\Lambda$-module $X$ such that $P^{\prime}\cong X/XJ$.
Proof. Assume that (1) and (2) are satisfied, and let $P$ be a relatively big projective right $\Lambda$-module. Then there exists a countably generated projective right $\Lambda$-module $Q$ with trace ideal $J$ such that $P$ is relatively $J$-big.
By (1), $Q$ can be taken to be a direct sum of finitely generated projective modules. By (2) applied to $P/PJ$, there exists a finitely generated projective right $\Lambda$-module $X$ such that $X/XJ\cong P/PJ$. Let $P^{\prime}=X\oplus Q^{(\omega)}$. By Lemma 1.17, $P^{\prime}$ is relatively big, and by Lemma 1.14, $P\cong P^{\prime}$, so it is a direct sum of finitely generated projective modules.
Assume now that every relatively big projective right $\Lambda$-module is a direct sum of finitely generated projective modules. Let $P$ be a projective right $\Lambda$-module with trace $J$. By [kaplansky, Theorem 1], $P=\bigoplus_{i\in I}P_{i}$ is a direct sum of countably generated projective modules. By Example 1.16, $P_{i}^{(\omega)}$ is relatively big. Therefore, $P_{i}^{(\omega)}$ is isomorphic to a direct sum of finitely generated projective modules for each $i\in I$. Hence, $P^{(\omega)}$ is isomorphic to a direct sum of finitely generated projective modules and $J=\operatorname{Tr}_{\Lambda}(P^{(\omega)})$. This gives (1).
To prove (2), let $Q$ be a countably generated projective right $\Lambda$-module with trace $J$, and let $P^{\prime}$ be a finitely generated projective right $\Lambda/J$-module. By Lemma 1.9, since $J$ is the trace of a countably generated projective module, there exists a countably generated projective right $\Lambda$-module $P$ such that $J\subseteq\operatorname{Tr}_{\Lambda}(P)$, $P/PJ\cong P^{\prime}$, and $\operatorname{Tr}_{\Lambda}(P)/J=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
By Lemma 1.17, $P\oplus Q^{(\omega)}$ is relatively big, so it is a direct sum of finitely generated projective modules. Write $P\oplus Q^{(\omega)}=\bigoplus_{i\in I}X_{i}$ with $X_{i}$ a finitely generated projective $\Lambda$-module for each $i\in I$. Since $(\bigoplus_{i\in I}X_{i})/(\bigoplus_{i\in I}X_{i})J\cong P^{\prime}$ is finitely generated, there is a finite subset $I_{0}\subseteq I$ such that $(\bigoplus_{i\in I_{0}}X_{i})/(\bigoplus_{i\in I_{0}}X_{i})J\cong P^{\prime}$, so $X=\bigoplus_{i\in I_{0}}X_{i}$ is the finitely generated projective $\Lambda$-module we were looking for.
1.2. Finite generation modulo an ideal and the ()(*)-condition
In this subsection, we take a closer look at condition (2) in Definition 1.15.
@@ 1.2. Finite generation modulo an ideal and the ()(*)-condition
We recall the following crucial properties of a descending sequence of right ideals associated with a column finite idempotent matrix. The result is taken from [traces], but all the necessary ideas are already in [P2].
@@
[traces, Lemmas 5.1 and 5.2] Let $\Lambda$ be a ring. Let $A=(a_{i,j})$ be an $\mathbb{N}\times\mathbb{N}$-matrix, with entries in $\Lambda$, that is a column finite matrix and such that $A^{2}=A$. Set $P=A\Lambda^{(\mathbb{N})}$. For any $k\geq 0$, let $I_{k}=\sum_{i>k,j\in\mathbb{N}}a_{ij}\Lambda$. Then,
@@
    $\mathcal{I}(P)$ has minimal elements if and only if there exists $k_{0}\in{\mathbb{N}}_{0}$ such that $\Lambda I_{k_{0}}=\Lambda I_{k_{0}+\ell}$ for every $\ell\in\mathbb{N}$. In this case, $I_{k_{0}}=I_{k_{0}}^{2}$, and $\Lambda I_{k_{0}}$ is the minimal element of $\mathcal{I}(P)$.
@@
[P2] A ring $\Lambda$ satisfies condition ()(*) provided that every descending chain of two-sided ideals
@@
The following statement explains the role of condition ()(*) in our discussion.
@@
Let $\Lambda$ be a ring. Let $P$ be a countably generated projective right $\Lambda$-module. If $\Lambda$ satisfies condition ()(*), then $\mathcal{I}(P)$ has a minimal element, and this minimal element is an idempotent ideal.
Proof. By Lemma 1.20 $(v)$, $\mathcal{I}(P)$ has a minimal element if and only if the descending sequence of two-sided ideals $(\Lambda I_{k})_{k\geq 0}$ is stationary. By Lemma 1.20 $(iii)$, this will happen if and only if the partial subsequence $(J_{n})_{n\geq 0}$ given by that statement is stationary. Since such a partial subsequence satisfies that $J_{n+1}J_{n}=J_{n+1}$, for any $n\geq 0$, it is stationary because $\Lambda$ satisfies condition ()(*).
@@
[P2, Lemma 2.8] Let $\Lambda$ be a noetherian ring. Let $P$ be a countably generated projective right $\Lambda$-module such that $\mathcal{I}(P)$ has a minimal element $J$. Then $J$ is the trace ideal of a countably generated projective right $\Lambda$-module, and $P$ is $J$-big. In particular, if $\Lambda$ satisfies condition ()(*), then any countably generated projective right module is relatively big.
@@
$(2)$ We do not know whether condition ()(*) is left-right symmetric, even for the case of noetherian rings.
$(3)$ There are noetherian rings with infinite descending sequences of idempotent ideals, so they do not satisfy condition ()(*). Over such rings there may be countably generated projective modules that are not relatively big. See, for example, the last section of [traces].
@@ 2. The monoid of relatively big projective modules
Let $\Lambda$ be a ring. We denote by $B(\Lambda)$ the subset of $V^{*}(\Lambda)$ of isomorphism classes of relatively big projective right $\Lambda$-modules which are not finitely generated. We denote by $V(\Lambda)\sqcup B(\Lambda)$ the disjoint union of $V(\Lambda)$ and $B(\Lambda)$, which is also a subset of $V^{*}(\Lambda)$, and consists of the isomorphism classes of all the relatively big projective right $\Lambda$-modules.
It is important to observe that, by Lemma 1.14, if $P$ is a countably generated projective right $\Lambda$-module such that $P]\in B(\Lambda)$, then $P$ is relatively big, so there exists a non-zero ideal $J\in\mathcal{T}(\Lambda)$ such that $P/PJ$ is finitely generated. Moreover, $J$ and $P/PJ$ uniquely determine $[P]$. We will use this fact freely throughout the paper.
@@
Proof. $(i)$. Let $P_{1}],[P_{2}]\in B(\Lambda)$. Then, for each $i=1,2$, there exists a countably generated non-zero projective right $\Lambda$-module $Q_{i}$ with trace ideal $J_{i}$ such that $P_{i}\oplus Q_{i}^{(\omega)}\cong P_{i}$ and $P_{i}/P_{i}J_{i}$ is a finitely generated projective right $\Lambda/J_{i}$-module. Take $P=P_{1}\oplus P_{2}$ and $Q=Q_{1}\oplus Q_{2}$. Then
@@
By Lemma 1.5, $J=\operatorname{Tr}_{\Lambda}(Q)=\operatorname{Tr}_{\Lambda}(Q_{1})+\operatorname{Tr}_{\Lambda}(Q_{2})=J_{1}+J_{2}$, so
@@
and $P/PJ$ is a finitely generated projective right $\Lambda/J$-module because it is a direct sum of quotients of finitely generated projective modules. Therefore, as $P$ is not finitely generated, $P]\in B(\Lambda)$.
Statement $(ii)$ is proved similarly, taking into account that the modules in $V(\Lambda)$ are the relatively 0-big projective modules.
@@
Let $\Lambda$ be a ring, and let $I\in\mathcal{T}(\Lambda)$. Then, there is a semigroup embedding $\iota_{I}\colon V(\Lambda/I)\hookrightarrow V(\Lambda)\sqcup B(\Lambda)$. The image of $\iota_{I}$ is all the isomorphism classes of relatively $I$-big projective right $\Lambda$-modules.
Proof. Let $Q$ be a countably generated projective right $\Lambda$-module with trace ideal $I$. Let $P^{\prime}$ be a finitely generated projective right $\Lambda/I$-module. By Lemma 1.9, there exists a projective right $\Lambda$-module $P$ such that $I\subseteq\operatorname{Tr}_{\Lambda}(P)$, $P/PI\cong P^{\prime}$, and $\operatorname{Tr}_{\Lambda}(P)/I=\operatorname{Tr}_{\Lambda/I}(P^{\prime})$.
By Lemma 1.17, $P\oplus Q^{(\omega)}$ is relatively $I$-big, and then $P\oplus Q^{(\omega)}]\in V(\Lambda)\sqcup B(\Lambda)$. The isomorphism class of $P\oplus Q^{(\omega)}$ is uniquely determined by $I$ and $P^{\prime}$. Hence, there is a well-defined embedding $\iota_{I}\colon V(\Lambda/I)\longrightarrow V(\Lambda)\sqcup B(\Lambda)$ defined by $\iota_{I}([P^{\prime}])=[P\oplus Q^{(\omega)}]$.
The map $\iota_{I}$ is a semigroup homomorphism because, for $P^{\prime}_{1}$ and $P^{\prime}_{2}$ in $V(\Lambda/I)$:
$\iota_{I}([P_{1}^{\prime}\oplus P_{2}^{\prime}])$ $=[P_{1}\oplus P_{2}\oplus Q^{(\omega)}]=[P_{1}\oplus Q^{(\omega)}\oplus P_{2}\oplus Q^{(\omega)}]$
$=[P_{1}\oplus Q^{(\omega)}]+[P_{2}\oplus Q^{(\omega)}]=\iota_{I}([P_{1}^{\prime}])+\iota_{I}([P_{2}^{\prime}]).$
where for $i=1,2$, $P_{i}$ denotes a countably generated projective module such that $P_{i}/P_{i}\cong P^{\prime}_{i}$, which exists because of Lemma 1.9.
Let $P$ be a relatively $I$-big projective right $\Lambda$-module. Then $P/PI$ is a finitely generated projective right $\Lambda/I$-module, and $\iota_{I}([P/PI])=[P]$.
If $I\neq 0$, the embedding $\iota_{I}$ of Lemma 2.3 is not a monoid embedding because $[0]$ maps to the isomorphism class of a non-finitely generated projective module.
@@
Consider the set $\mathcal{T}(\Lambda)$ and the disjoint union $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$. We will denote the elements of $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$ by $([P],I)$ with $P]\in V(\Lambda/I)$. Note that if $I\subseteq J$, there is a morphism of monoids $f_{JI}\colon V(\Lambda/I)\to V(\Lambda/J)$ defined by $f_{JI}([X],I)=([X/XJ],J)$ with $X]\in V(\Lambda/I)$. Therefore, $f_{JI}(V(\Lambda/I))\subseteq V(\Lambda/J)$. Define addition in $\bigsqcup_{I\in\mathcal{T}(\Lambda)}V(\Lambda/I)$ by
$([P_{1}],I)+([P_{2}],J)$ $=f_{I+J,I}([P_{1}],I)+f_{I+J,J}([P_{2}],J)$
@@
for each $P_{1}]\in V(\Lambda/I)$, $P_{2}]\in V(\Lambda/J)$. Therefore, the universal property of the disjoint union and Lemma 2.3 imply that there exists a monoid embedding
@@
The following result states that $\iota$ is an isomorphism of monoids.